OpenStudy (kenialove):

Earths atmosphere is 78% nitrogen by volume. About how much nitrogen is in a tank containing 600 cubic meters of air ? A. 132 m^3 B. 169 m^3 C. 420 m^3 D. 468 m^3

OpenStudy (michele_laino):

the requested volume is given, by definition of percentage, by this computation: \[\Large V = \frac{{78}}{{100}} \cdot {V_{TOTAL}} = \frac{{78}}{{100}} \cdot 600 = ...?\]

OpenStudy (kenialove):

So the answer is D. 468 m^3 ? 78/100 = 0.78 0.78 x 600 = 468

OpenStudy (michele_laino):

yes!

OpenStudy (kenialove):

Thanks so much

OpenStudy (michele_laino):

:)